MARS DESERT RESEARCH STATION

Green Hab Report – February 3rd

Green Hab Report:
Prepared by: Sneha Velayudhan
Sol 05
Earth Date: 02.03.2017

Functionality: A sunny day on Mars means regular monitoring of the Green Hab as temperature rise quickly. Morning the temperatures were measured. The outside temperature was 9 deg C while the green hab temperature was 38 deg C and under misian Mars lamp it was 20 deg C at around 10:30 AM. Post lunch, the skies were clear and the temperature was high. It was an indication to me that the temperatures in the Green Hab would be high. I went to the green hab to find the inside temperature to be nearly 50 deg C. I switched on the cooler and watered the plants again. The humidity was checked and it was around 30 %. The cooler was left on in order to control the temperature of the Green Hab. In the evening, the outside temperature was 8 deg C, Green Hab temperature was 20 deg C and under the misian Mars Lamp temperature was 22.5 deg C. The Humidity noted inside the hab was 15% and under the misian Mars lamp was 30%.

Status: In the green hab, the plants and seedlings were watered. Special attention given to the Fenugreek seeds. The root growth in the Earth Soil was impressive. The leaves were sprouting out of the seed coat and it was greenish. It was a great relief to see the seedlings grow!!! Unfortunately, there is no much growth observed in the seeds sowed in the Mars Soil. It was expected but that was little depressing 🙁

Under the misian Mars Lamp, the seedlings were watered and observed for growth. The earth soil seeds showed some growth but it looks like plants enjoy sunshine more that the artificial lamp light! Similar to the green hab, the Mars soil pots did not show any impressive growth. The root growth was measured in both Earth and Mars seedlings. In the evening the plants were watered with Vit D in incremental measurement and the temperature was monitored.

Planned activity: Monitoring and maintaining the Green Hab temperature. Tabulating the growth of plants with varied amount of Vitamin D.

Sol Summary – February 2nd

Sol Summary-04
Earth Date: 02.02.2017
Finally the day was here when our Commander Mamatha joined us 🙂
should mention it as the most awaited day of our mission!! After a
long night of discussions about how to pick her, how to communicate
with her, who should go and when to go, it was decided that Saroj and
Arpan would go to pick her while Sneha and Arun will stay in the hab
performing the experiments. Saroj known as our official driver, not to
forget the second in command was planning the activities to be done
while they were away.
It must have been a long night for Arun as he worked all night reading
materials for his geological research. As morning came in, the crew
woke up as usual and started with the normal duties. Water supply to
the hab was still and issue, water was filled in vessels and stored
for use. Last night we had received instructions on how to clean the
filter, Saroj was trying hard to download and read them, unfortunately
he was not able to download all the steps as the internet was very
slow this morning. Three crew members had breakfast together while the
fourth was asleep. After breakfast and some discussion about the day,
some music, jokes and fun. It was a fun filled day as we were excited
about the arrival of Commander. Sneha went to the green hab watered
the plants spent time doing her experiment. Saroj and Arpan prepared
lunch. It was an early lunch as they had to leave to Green River.
After the operations check, the reports were prepared. After lunch
Saroj and Arpan got ready to leave to pick up Mamatha. The second in
command was so excited that it was decided the two men would follow
the protocol and wear jump suits to receive the commander !!! Sneha
did comment that when the Commander sees this, she would also love to
wear the jump suit and join the team before driving to Mars 🙂
After they left, it was a very long day on Mars, Arun was reading his
materials and noting the places for his next EVA. Sneha remained in
the hab and was trying to make some bread for the next few days. The
bread turned out to be disastrous 🙁 The long day came to an end when
Mamatha arrived on Mars. There is excitement, screams of joy, reunion
of friends after 2 years. Its really exciting now.

Green Hab Report – February 2nd

Green Hab report


Written by : Sneha Velayudhan
Sol 04
Earth Date: 02.02.2017
Functionality: Its a nice sunny day on Mars. Temperature at around
10:20 AM was 2 deg. C outside, while the green hab temperature was 22
deg. C and the misian Mars lamp temperature was 19 deg. C. The
temperature was again checked in the evening at 5:30 PM, inside the
greenhab the temperature was 30 deg. C, it was around 22.4 deg. C
under the misian Mars lamp while the outside temperature was 9 deg. C
and windy.
Status: The temperature was monitored and seedlings were watered.
Special attention was given to the fenugreek seeds sowed yesterday as
a part of botany experiment. It was found that the soil in the pots
was mostly dry. In order to maintain the adequate moisture in the
soil, the pots were watered in the green hab and under the misian Mars
lamp. Some amount of water was poured into the tray beneath the pots
to maintain regular moisture.
By the evening the temperature was monitored again and the seedlings
were watered. The fenugreek was watered with fixed amount of water and
Vit. D ranging from 0.1 gm to 1 gm both in the green hab and under the
misian Mars lamp. The root growth was measured and tabulated
accordingly.
Planned activity: The seedlings will be monitored for growth along
with the temperature and humidity.

Crew Photos – February 1st

Arpan (Engineer) and SaiArun (Geologist) in the airlock prior to their EVA
Arpan and Sneha posing casually on a bright sunny Mars day
Crew Biologist (Sneha) filling the Earth and Mars soil to test plant growth
Crew Engineer (Arpan) and Crew Scientist (Saroj) trying to figure out issue with overhead water system
Crew Engineer and Geologist on their way to explore Mars
Crew enjoying delicious lunch after a tiring EVA
Preperation of ‘Mars soil’ to moniter plant growth

Sol Summary – February 1st

Journalist report Sol 03

It has been just corn flakes for the breakfast ever since we
landed on Mars, so we started our day with pasta for a change which
happens to be one of my favourite dishes! I appreciate the extra
efforts Italians put to get the perfect recipes and how particular
they are with the cooking. I recollect our commander Mamatha’s
experience with her Italian friend who accidentally spotted the over
cooked pasta during lunch break and presented mamatha a timer clock
the next day with a tag just 10 minutes!!! The breakfast discussion
was about our geologist Arun and scientist Saroj’s previous night star
trail photo session inside tunnel which turns out to be hilarious! We
had a laughter riot when Saroj recalled how Arun was scared of Martian
cats attack!! Yes we heard stories of the Martian cats spotting from a
very close source 😉 They did also spot a beam of light near the solar
panel area, it must be of an alien spaceship!! Amidst chaos and in the
verge of getting to the hab quickly, Arun had his heart in the mouth
after dropping his camera. He was relieved after finding out just the
UV filters was shattered!
Post breakfast crew biologist Sneha sowed the germinated
seeds in pots. She will be studying the effects of Vitamin D on the
growth of plants. Each set of pots have been placed in misian Martian
lamp and greenhab area. By noon, Arun and Saroj left for the planned
EVA, the primary goal was to collect rock samples with mineral
composition to help Arun with his geology research. We were expecting
quite a lot of pictures of the activity but sadly it turns out the SD
card was forgotten!!! I guess there is a camera curse and the team
intends to remove the same using age old Indian technique “kala teeka”
(removal of bad omens).
Sol 04 is going to be an exciting day, as our commander will be
joining us. Can’t wait to start the planetary protection studies.

Green Hab Report – February 1st

Green Hab Report
Sol-03
Earth Date: 02.01.2017
Written by : Sneha Velayudhan

Functionality: as the climate on Mars is beautiful, the green hab
temperatures are high that we had to switch on the coolers twice
today. The temperatures were noted as 28 degree centigrade while the
outside temperature was 1 degree centigrade at 10:10 AM. By 12:20 pm,
the temperature rised to around 45 degree, the cooler was switched on
to decrease the temperature. The temperature was brought down to
around 30 degree. In the evening temperatures were noted, the green
hab was 18 degree while the outside temperature was 9 degree
centigrade.
Status: The botany experiment started with 2 sets of pots, 10 each
prepared with earth soil and marked with 0-0.9 grams of Vitamin D.
Five different types of Martian soil, each type of soil were prepared
in 2 pots and marked for the variety. A mixture of earth soil, the
different tryps of Martian soil was prepared in equal ratio and filled
into 4 pots. Each of the respective sets such as the 10 pots of Earth
soil, 5 pots of mars soil, 2 pots each of the mixture of Earth and
Mars soil were consolidated. The germinated fenugreek seeds were
delicately sowed into the pots.  One set is placed in the green hab
and the other under the misian mars lamp. All plants including the
newly potted seedlings were watered twice today.
Planned activity: The growth of fenugreek will be monitored along with
the other plants in the coming days.

EVA Report – February 1st

EVA Report:
SOL: 03
Date: 02/01/2017
Person filling out the report: Sai Arun Dharmik
Crew members involved in the EVA: Sai Arun Dharmik and Saroj Kumar
EVA leader: Sai Arun Dharmik
Begin: 11:45 Hrs
End: 15:15 Hrs
Type of EVA: Deimos + 1 ATV (300) + Walking
Purpose: Collection of rock samples to be measured magnetic susceptibility
UTM Coordinates: Northing: 4255600 , Easting : 518400
Summary: We started the EVA at 11:45 am. The temperature was around 12
degree Celsius outside the Hab and the sky was clear. We have
collected few samples from the URC North site area on our route to
Lith Canyon. The samples collected at URC North site were primarily
transported due to weathering, hence we cannot attribute them to that
location. Later we collected few samples of sandstones and shales near
the Lith canyon region. The samples will be analysed tomorrow
(02.02.2017) and a detailed report will be sent.

Crew Photos – January 31st

Crew members studying geology map of Mars.

 

Crew biologist Sneha potting the Seedlings into soil with help from crew engineer.

 

Crew biologist digging soil with shovel.

 

Crew biologist and crew engineer in the field.

 

Crew biologist and crew engineer ready for EVA.

 

Sprouted fenugreek seedlings.

Sol Summary – January 31st

Crew 174 Daily Summary Sol – 02
We started our day with a positive news of our Crew Commander’s arrival! After the breakfast the crew members geared up for the planned activities. The first EVA experience came handy today, the preparation time of wearing EVA suits and checking radio communications was short. As per our planned EVA, the crew biologist and scientist quickly left for the same. The geological map of Mars was studied in prior to mark locations for soil collection. The EVA lasted around 2 hours and 30 mins, during the EVA the crew was able to look for 5 different types of soil in color and texture. They captured photos and videos of the red planet. The crew returned to the hab after the regular hab operations checks. In parallel, crew in the hab prepared lunch and communicated with the EVA team when needed. After EVA return, the crew members had lunch and rested.
Later, Arpan and Saroj helped Sneha organize the soil collection and prepare the pots to plant the seeds with the Martian and Earth soil with added Vitamin D, to observe their growth patterns in controlled conditions. After the soil prep, the team started to work on the reports and planned activities for tomorrow.

GreenHab Report – January 31st

Green Hab Report
Report written by: Sneha Velayudhan (Crew Biologist)
Date : 01/31/2017
Functionality: The heaters in the green hab is in perfect working condition. The temperatures are monitored twice a day. Today it was noted that the temperature in the green hab was around 48 degree centigrade while the temperature outside was 10 degree centigrade around 11:30 AM. As the temperature was too high the cooler was switched on and the green hab temperature was brought down to around 33 degree centigrade. The seedlings were watered twice today. The temperatures were monitored in the evening, it was noted to be 23 degree centigrade inside while the outside temperature was around 8 degree centigrade.
Status: The fenugreek and the Chick pea seeds have germinated well and are ready for potting.  The pots are prepared with earth soil for the experiment to monitor the effect of Vitamin D on the growth of plants. An EVA was done to collect different types of soil from different locations on Mars. There were 5 different types of soil collected. The soil collected are different in color and texture. Seeds will be planted in each of the soil type and the seedling growth will be observed. The existing seedlings were monitored for growth. The height of the corn plants were measured.
Planned activity: The germinated seeds will be planted into the earth and martian soil. There will be 2 sets each of 10 pots marked in increments of 0 – 0.9 g of Vitamin D. One of this set will be placed in the Green Hab and the other under the misian Mars light and growth will be monitored.